OpenStudy (anonymous):

A motorboat takes 5 hours to travel 200 miles going upstream. The return trip takes 4 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (jdoe0001):

\(\large{ b=\textit{boat's rate}\qquad s=\textit{stream rate} \\ \quad \\ \begin{array}{rcccllll} &distance&rate&time \\\hline\\ upstream&200&b-r&5\\ downstream&200&b+r&4 \end{array}\implies \begin{array}{llll} {\color{brown}{ 200}}=(b-r)5 \\ \quad \\ {\color{brown}{ 200}}=(b+r)4 \end{array} \\ \quad \\ \implies (b-r)5=(b+r)4}\) solve for "b" or "r" then grab that value and substitute on either of the two equations above to get the other variable